THE Ã(0, 8, 0)^{1} ← X̃(0, 0, 0)^{2} AND Ã(0, 8, 0)^{3} ← X̃(0, 0, 0)^{2} BANDS OF CH_{2}^{+}

\underline{CHRISTOPHER F. NEESE}, JENNIFER L. GOTTFRIED, and TAKESHI OKA, Department of Chemistry, Department of Astronomy & Astrophysics, and the Enrico Fermi Institute, The University of Chicago, Chicago IL, 60637, USA.

The methylene ion, CH_{2}^{+}, is of special theoretical interest because it is both quasi-linear and exhibits a strong Renner-Teller interaction between its ground and first-excited electronic states. At linearity, the ground state is a ^{2}Π_{u} state that splits into X̃^{2}A_{1} and Ã^{2}B_{2} states as the molecule bends. The Ã state is linear, while the X̃ state is quasi-linear with a barrier to linearity of only 1089 cm^{−1}.

Recently, the spectrum of CH_{2}^{+} in the region 10,700–11,000 cm^{−1} has been recorded with our Ti:sapphire laser spectrometer. This spectrometer couples velocity modulation with heterodyne detection for near shot-noise-limited sensitivity. In this new spectrum we have assigned the Ã(0, 8, 0)^{1} ← X̃(0, 0, 0)^{2} and Ã(0, 8, 0)^{3} ← X̃(0, 0, 0)^{2} bands.^{a} These bands are the first high-resolution detection of K_{a} = 2 levels for the ground state. In addition, the Ã(0, 8, 0)^{1} ← X̃(0, 0, 0)^{2} can be combined with the previously studied Ã(0, 8, 0)^{1} ← X̃(0, 0, 0)^{0} band^{b} to produce K_{a} = 2 − 0 combination differences for the ground state. These combination differences provide the first high-resolution information about the ground state A rotational constant.

aThe X̃ state is labeled using bent notation and the Ã state is labeled using linear notation.

bJ. L. Gottfried and T. Oka, J. Chem. Phys. 121, 11527 (2004).